The following facts pertain to a noncancelable lease agreement between Sandhill Leasing Company and Teal Company, a lessee.

Inception date: May 1, 2017
Annual lease payment due at the beginning of
   each year, beginning with May 1, 2017 $19,373.99
Bargain-purchase option price at end of lease term $4,400
Lease term 5 years
Economic life of leased equipment 10 years
Lessor’s cost $62,000
Fair value of asset at May 1, 2017 $85,000
Lessor’s implicit rate 9 %
Lessee’s incremental borrowing rate 9 %


The collectibility of the lease payments is reasonably predictable, and there are no important uncertainties surrounding the costs yet to be incurred by the lessor. The lessee assumes responsibility for all executory costs. The expected residual value of the equipment at the end of 5 (10) years is $12,000 ($0).
Answer #1
Teal Company (Lessee)
Lease Amortisation Schedule
Date Annual Lease Payment + BPO Interest on liabilities Reduction in lease liability Lease Liability
(A) (B=D×9%) (C=A-B) (D)
5/1/17 $85,000
5/1/17 $19,373.99 - $19,373.99 $65,626.01
5/1/18 $19,373.99 $5,906.34 $13,467.65 $52,158.36
5/1/19 $19,373.99 $4,694.25 $14,679.74 $37,478.62
5/1/20 $19,373.99 $3,373.08 $16,000.91 $21,477.71
5/1/21 $19,373.99 $1,932.99 $17,441.00 $4,036.71
4/30/22 $4,400 $363.30 $4,036.70 -